# Homological shifts of a complementary edge ideal

### v1

@@ 4.1. The case when $i=1$
As proven in [13], for polymatroidal ideals $I$, $\mathrm{HS}_{1}(\mathcal{R}(I))$ is generated in degree $1$ with a relatively straightforward proof. The case $i=1$ for complementary edge ideals also admits a similarly elementary proof.
@@
Let $G$ be a connected graph on the set $[n]$ and let $I=I_{c}(G)$ be its complementary edge ideal. Then $\mathrm{HS}_{1}(\mathcal{R}(I))$ is generated in degree $1$.
@@
It suffices to prove that $\mathrm{HS}_{1}(I^{s+1})=I^{s}\cdot\mathrm{HS}_{1}(I)$ for all $s\geq 0$. Let $\alpha=\prod\limits_{i=1}^{n}x_{i}$. It is easy to see that $\mathrm{HS}_{1}(I)=\left(\frac{\alpha}{x_{i}}\mid\deg_{G}(i)\geq 2\right)$. We will prove the statement by induction on $s$, and the case where $s=0$ is trivial. Suppose that $s\geq 1$. For any $\beta\in\mathcal{G}(\mathrm{HS}_{1}(I^{s+1}))$. Then, by [18, Lemma 2.3], there are two elements $u,v\in\mathcal{G}(I^{s+1})$ such that $\beta=\frac{\alpha^{s+1}}{\gcd(u,v)}$, where $\deg(\gcd(u,v))=2s-1$. We can write $u$ as $u=\prod\limits_{i=1}^{s+1}u_{i}$ where each $u_{i}\in\mathcal{G}(I(G))$. Therefore, $u=y_{1}\gcd(u,v)$. Without loss of generality, we can assume that $y_{1}$ divides $u_{1}$ and that $u_{1}=x_{i}y_{1}$ for some $i\in[n]$. Then $\gcd(u,v)=x_{i}\prod\limits_{i=2}^{s+1}u_{i}$. We consider the following two cases:
(1) If $\deg_{G}(i)\geq 2$, then $\beta$ can be written as
$\beta=\frac{\alpha^{s}}{u_{2}\cdots u_{s+1}}\cdot\frac{\alpha}{x_{1}},$
and the conclusion follows.
(2) If $\deg_{G}(i)=1$, then $v$ can be written as $v=\prod\limits_{i=1}^{s+1}v_{i}$. Assuming that $x_{i}$ divides $v_{1}$, we have $u_{1}=v_{1}$. In this case, we set $f=\prod\limits_{i=2}^{s+1}u_{i}$ and $g=\prod\limits_{i=2}^{s+1}v_{i}$. Then $\gcd(u,v)=u_{1}\gcd(f,g)$. Therefore, $\frac{\alpha^{s}}{\gcd(f,g)}\in\mathcal{G}(\mathrm{HS}_{1}(I^{s}))$. By the induction hypothesis, $\frac{\alpha^{s}}{\gcd(f,g)}\in\mathcal{G}(I^{s-1}\cdot\mathrm{HS}_{1}(I))$. Therefore, $\beta=\frac{\alpha}{u_{1}}\cdot\frac{\alpha^{s}}{\gcd(f,g)}\in\mathcal{G}(I^{s}\cdot\mathrm{HS}_{1}(I))$, as desired.    ∎
@@ 4.2. Trees
Let $G$ be a tree with $n$ vertices, labelled as in Remark 3.5, and let $I=I_{c}(G)$ denote its complementary edge ideal. Define a map $\phi:[n-1]\rightarrow[n]$ such that $\phi(k)$ is the unique $j\in[n]$, with $j>k$ for which $\{k,j\}\in E(G)$. Then, for any $1\leq i\leq s$, we have
$\mathrm{HS}_{i}(I^{s})=I^{s-i}\cdot\left(\frac{\alpha^{i}}{\prod_{k\in F}x_{\phi(k)}}:\ F\subseteq[n-2]\text{\ and\ }|F|=i\right).$
@@
Let $e_{1},e_{2},\ldots,e_{n-2},e_{n-1}$ be all the edges of $G$. Without loss of generality, we can assume that $\min(e_{i})=i$ for all $i\in[n-1]$. Thus, $e_{i}=x_{i}x_{\phi(i)}$. Note that every element of $\mathcal{G}(I(G)^{s})$ is uniquely written as
@@
where $\sum\limits_{i=1}^{n-1}s_{i}=s$ and each $s_{i}\geq 0$. Furthermore, $e_{1},\ldots,e_{n-1}$ are algebraically independent.
For any $\beta=\frac{\alpha^{s-i}}{e_{1}^{a_{1}}\cdots e_{n-1}^{a_{n-1}}}\in\mathcal{G}(I^{s-i})$ and for any $F\subseteq[n-2]$ with $|F|=i$, we define
@@
where $b_{i}=1$ if $i\in F$, and $b_{i}=0$ otherwise. In particular, $b_{n-1}=0$, so $\gamma\in\mathcal{G}(I^{s})$, and $F\subseteq\operatorname{set}(\gamma)$ by the equation (3). This implies that ${\bf x}_{F}\gamma\in\mathrm{HS}_{i}(I^{s})$. Note that
${\bf x}_{F}\gamma=\beta\frac{\alpha^{i}}{\prod_{k\in F}x_{\phi(k)}},$
we conclude that the inclusion $\supseteq$ holds.
Conversely, for any $U\in\mathcal{G}(\mathrm{HS}_{i}(I^{s}))$, we can write $U$ as
$U={\bf x}_{F}V={\bf x}_{F}\frac{\alpha^{s}}{e_{1}^{s_{1}}\cdots e_{n-1}^{s_{n-1}}},$
where $V=\frac{\alpha^{s}}{e_{1}^{s_{1}}\cdots e_{n-1}^{s_{n-1}}}\in\mathcal{G}(I^{s})$ and $F\subseteq\operatorname{set}(V)$ with $|F|=i$. By the equation (3), we see that $s_{i}>0$ for all $i\in F$. Let $c_{i}=s_{i}-1$ if $i\in F$, and $c_{i}=s_{i}$ otherwise. Then $W:=\frac{\alpha^{s-i}}{e_{1}^{c_{1}}\cdots e_{n-1}^{c_{n-1}}}\in\mathcal{G}(I^{s-i})$. Furthermore, we have $W\frac{\alpha^{i}}{\prod_{i\in F}x_{\phi(i)}}={\bf x}_{F}V$. This implies that the inclusion $\subseteq$ holds, completing the proof.    ∎
@@
Let $G$ be a tree and let $I=I_{c}(G)$ be its complementary edge ideal. Then the $i$-th homological shift algebra $\mathrm{HS}_{i}(\mathcal{R}(I))$ is generated in degree $i$ for $1\leq i\leq n-2$.
@@
By Theorem 4.3, it follows that $\mathrm{HS}_{i}(I^{s})=I^{s-i}\cdot\mathrm{HS}_{i}(I^{i})$ for all $s\geq i$. This implies $\mathrm{HS}_{i}(\mathcal{R}(I))$ is generated in degree $\leq i$. But $\mathrm{HS}_{i}(I^{s})=0$ for $s<i$, the result follows.    ∎
The ideal $\left(\frac{\alpha^{i}}{\prod_{k\in F}x_{\phi(k)}}:\ F\subseteq[n-2]\text{\ and\ }|F|=i\right)$ is uniquely determined by $G$ and $i$ up to the permutation of variables, since it equals to $\mathrm{HS}_{i}(I^{i})$. This implies it is independent of the labelling of $G$. We denote this ideal by $J_{i}(G)$.
To study the properties of $J_{i}(G)$, we recall some theory on polymatroidal ideals. We follow the notation from [23] and [6]. Given a vector ${\bf a}=(a_{1},\dots,a_{n})\in{\mathbb{Z}}_{\geq 0}^{n}$ and an integer $d$ with $a_{1}+\dots+a_{n}\geq d$, the ideal of Veronese type (determined by $\bf a$} and $d$) is the polymatroidal ideal $I_{{\bf a},d}$ with minimal monomial generating set
$\mathcal{G}(I_{{\bf a},d})\ =\ \{{\bf x^{b}}\in S:\ |{\bf b}|=d,\ {\bf b}\leq{\bf a}\}.$
Here, if ${\bf b}=(b_{1},\dots,b_{n})\in{\mathbb{Z}}_{\geq 0}^{n}$ we put ${\bf b}(i)=b_{i}$, $|{\bf b}|=\sum_{i=1}^{n}{\bf b}(i)$ is the modulus of $\bf b$} and ${\bf b}\leq{\bf a}$ means that ${\bf b}(i)\leq{\bf a}(i)$ for all $i$.
A monomial ideal $I\subset S$ generated in a single degree is said to satisfy the strong exchange property if: for all $u,v\in\mathcal{G}(I)$, all $i$ such that $\deg_{x_{i}}(u)>\deg_{x_{i}}(v)$ and all $j$ such that $\deg_{x_{j}}(u)<\deg_{x_{j}}(v)$ we have $x_{j}(u/x_{i})\in I$. It is clear that an ideal of Veronese type satisfies the strong exchange property. By the work [24], it is known that ideals satisfying the strong exchange property are essentially ideals of Veronese type in the sense that if $I$ satisfies the strong exchange property, then $I=(u)I_{{\bf a},d}$ for some monomial $u\in S$, and some ${\bf a}\in{\mathbb{Z}}_{\geq 0}^{n}$ and $d>0$.
@@
Let $G$ be a tree with $n$ vertices. Then $J_{i}(G)$ is a poymatroidal ideal with the strong exchange property for all $i\in[n-2]$.
@@
Let $K_{i}(G)=\left(\prod_{k\in F}x_{\phi(k)}\mid F\subseteq[n-2]\text{\ and\ }|F|=i\right)$. For any $k\in[n-1]$, define
$b_{k}=|\phi^{-1}(k)|=\left|\{j\in[n-2]:\phi(j)=k\}\right|.$
@@
$K_{i}(G)=\left(\prod\limits_{k=1}^{n-1}x_{k}^{a_{k}}\mid\sum\limits_{k=1}^{n-1}a_{k}=i\text{\ and\ }0\leq a_{k}\leq b_{k}\ \text{for all}\ k\in[n-1]\right).$
For every $F\subseteq[n-2]$ such that $|F|=i$, let $a_{k}=|F\cap\phi^{-1}(k)|$ for any $k\in[n-1]$. Then $0\leq a_{k}\leq b_{k}$ for all $k$, and
@@
Note that $\prod_{k\in F}x_{\phi(k)}=\prod\limits_{k=1}^{n-1}x_{k}^{a_{k}}$, so the inclusion $\subseteq$ holds.
Conversely, take any monomial $\prod\limits_{k=1}^{n-1}x_{k}^{a_{k}}$ such that $\sum\limits_{k=1}^{n-1}a_{k}=i$ and $0\leq a_{k}\leq b_{k}$ for all $k\in[n-1]$. Let $F=F_{1}\sqcup\cdots\sqcup F_{n-1}$, where $F_{k}$ is a subset of $a_{k}$ elements in $\phi^{-1}(k)$, and by convention, $F_{k}=\emptyset$ if $a_{k}=0$. Then $\prod_{k\in F}x_{\phi(k)}=\prod\limits_{k=1}^{n-1}x_{k}^{a_{k}}$, so the inclusion $\supseteq$ holds. This proves the claim.
Therefore, $K_{i}(G)$ is the ideal $I_{{\bf b};i}$ of Veronese type on the ground set $\phi([n-2])$. Here, ${\bf b}=(b_{1},\ldots,b_{n-1}).$ By putting $b_{n}=0$, we conclude that
$J_{i}(G)$ $=\left(\prod\limits_{k=1}^{n}x_{k}^{i-a_{k}}\mid\sum\limits_{k=1}^{n-1}a_{k}=i\text{\ and\ }0\leq a_{k}\leq b_{k}\ \text{for all}\ k\in[n]\right)$
$=\left(\prod\limits_{k=1}^{n}x_{k}^{c_{k}}\mid\sum\limits_{k=1}^{n-1}c_{k}=(n-2)i\text{\ and\ }i-b_{k}\leq c_{k}\leq i\ \text{for all}\ k\in[n]\right)$
$=\left(\prod\limits_{k=1}^{n}x_{k}^{i-b_{k}}\right)\left(\prod\limits_{k=1}^{n-1}x_{k}^{c_{k}-i+b_{k}}\mid\sum\limits_{k=1}^{n-1}c_{k}=(n-2)i\text{\ and\ }i-b_{k}\leq c_{k}\leq i\ \text{for all}\ k\in[n]\right)$
$=\left(\prod\limits_{k=1}^{n}x_{k}^{i-b_{k}}\right)\left(\prod\limits_{k=1}^{n-1}x_{k}^{a_{k}}\mid\sum\limits_{k=1}^{n-1}a_{k}=n-2-i\text{\ and\ }0\leq a_{k}\leq b_{k}\ \text{for all}\ k\in[n]\right)$
@@
Here, the second to last equality holds because of $\sum\limits_{i=1}^{n-1}b_{i}=n-2$. In the last equality, we use the convention that $K_{0}(G)=S$. Therefore, $J_{i}(G)$ is a polymatroidal ideal with the strong exchange property.    ∎
In fact, $J_{i}(G)$ is the dual polymatoidal ideal of $K_{i}(G)$ with respect to the cage the $n$-tuple $(i,\ldots,i)$ in the following sense. Let $I$ be a polymatoidal ideal. A cage for $I$ is an integral vector ${\bf c}\in\mathbb{N}^{n}$ such that ${\bf u}(i)\leq{\bf c}(i)$ for any $\bf u$} with ${\bf x}^{\bf u}\in\mathcal{G}(I)$ and any $i\in[n]$. Fix a cage $\bf c$} of $I$. Then the ideal generated ${\bf x}^{\bf c-u}:\ {\bf u}\in G(I)$ is also a polymatroid ideal, which we call the dual polymatroidal ideal of $P$ with respect to $\bf c$}. For more details, see [6] and [13].
In Lemma 4.5, we prove that the $i$-th homological shift ideal $\mathrm{HS}_{i}(I_{c}(G)^{i})(=J_{i}(G))$, when dividing a suitable monomial, is a ideal of Veronese type. In fact that every Veronese-type ideal arises in this way. Recall that a caterpillar tree is a tree $T$ with the property that it contains a path $P$ such that any vertex of $T$ is either a vertex of $P$ or it is adjacent to a vertex of $P$. Caterpillar trees were first studied by Harary and Schwenk [21]. A leaf edge is an edge that connects a leaf vertex (a vertex of degree 1) to another vertex in the graph.
@@
Let $I=I_{\mathbf{a},d}\subseteq\mathbb{K}[x_{1},\ldots,x_{n}]$ be a Veronese-type ideal, where $\mathbf{a}=(a_{1},\ldots,a_{n})\in\mathbb{Z}_{>0}^{n}$ and $d\leq|\mathbf{a}|$ (with $|\mathbf{a}|=a_{1}+\cdots+a_{n}$). Then there exists a caterpillar tree $T$ such that
$\mathrm{HS}_{|\mathbf{a}|-d}\left(I_{c}(T)^{|\mathbf{a}|-d}\right)=uI$
for some monomial $u$.

### v5

@@ 5.1. The case when $i=1$
Recall from [15, Lemma 2.3] that if $I$ is a monomial ideal (regardless of whether it has linear quotients), then $\mathrm{HS}_{1}(I)$ is generated by all $\mathrm{lcm}(u,v)$, where $u,v\in\mathcal{G}(I)$ and $u\neq v$. Here, $\mathrm{lcm}(u,v)$ denotes the least common multiple of $u$ and $v$.
@@
Let $G$ be a connected graph on the set $[n]$ and let $I=I_{c}(G)$ be its complementary edge ideal. Then $\operatorname{HS}_{1}(\mathcal{R}(I))$ is generated in degree $1$ as an $\mathcal{R}(I)$-module.
@@
It suffices to prove that $\mathrm{HS}_{1}(I^{s+1})=I^{s}\cdot\mathrm{HS}_{1}(I)$ for all $s\geq 0$.
First, we recall two definitions for clarity:
- For any vertex $i\in V(G)$, $\deg_{G}(i)$ denotes the number of edges incident to $i$.
- For two monomials $u,v\in S$, $\gcd(u,v)$ stands for the greatest common divisor of $u$ and $v$.
By [15, Lemma 2.3] and the fact that $\mathrm{HS}_{1}(I)$ is generated in degree n1n-1, it is straightforward to verify that $\mathrm{HS}_{1}(I)=\left(\frac{\alpha}{x_{i}}\mid\deg_{G}(i)\geq 2\right)$. We now proceed by induction on $s$. Note that $I^{0}=S$, so the case when $s=0$ is trivial.
Suppose $s\geq 1$. Let $\beta\in\mathcal{G}(\mathrm{HS}_{1}(I^{s+1}))$. By [15, Lemma 2.3], there exist elements $u,v\in\mathcal{G}(I(G)^{s+1})$ such that $\beta=\mathrm{lcm}\left(\frac{\alpha^{s+1}}{u},\frac{\alpha^{s+1}}{v}\right)=\frac{\alpha^{s+1}}{\gcd(u,v)}$, where $\deg(\gcd(u,v))=2s+1$. It follows that $u=y\cdot\gcd(u,v)$ for some $y\in\{x_{1},\ldots,x_{n}\}$.
Express $u$ as $u=\prod_{j=1}^{s+1}u_{j}$, where each $u_{j}\in\mathcal{G}(I(G))$. Without loss of generality, assume $y$ divides $u_{1}$, so $u_{1}=x_{i}y$ for some $i\in[n]$. Thus, $\gcd(u,v)=x_{i}\cdot\prod_{j=2}^{s+1}u_{j}$. We consider two cases:
(1) If $\deg_{G}(i)\geq 2$, then
$\beta=\frac{\alpha^{s}}{u_{2}\cdots u_{s+1}}\cdot\frac{\alpha}{x_{i}}\in I^{s}\cdot\mathrm{HS}_{1}(I).$
(2) If $\deg_{G}(i)=1$, express $v$ as $v=\prod_{j=1}^{s+1}v_{j}$, where each $v_{j}\in\mathcal{G}(I(G))$. Since $x_{i}\mid\gcd(u,v)$, we may assume $x_{i}$ divides $v_{1}$, which implies $u_{1}=v_{1}$ (as both are generators of $I(G)$ containing $x_{i}$). Let $u_{1}=v_{1}=w$ for simplicity. Set $f=\prod_{j=2}^{s+1}u_{j}$ and $g=\prod_{j=2}^{s+1}v_{j}$. Then $\gcd(u,v)=w\cdot\gcd(f,g)$, and $\frac{\alpha^{s}}{\gcd(f,g)}\in\mathcal{G}(\mathrm{HS}_{1}(I^{s}))$. By the induction, it follows that $\frac{\alpha^{s}}{\gcd(f,g)}\in I^{s-1}\cdot\mathrm{HS}_{1}(I)$. Therefore:
$\beta=\frac{\alpha}{w}\cdot\frac{\alpha^{s}}{\gcd(f,g)}\in I^{s}\cdot\mathrm{HS}_{1}(I),$
as required.
@@ 5.2. Trees
Let $G$ be a tree with $n$ vertices, labelled as in Remark 4.5, and let $I=I_{c}(G)$ denote its complementary edge ideal. Define a map $\phi:[n-1]\rightarrow[n]$ such that $\phi(k)$ is the unique $j\in[n]$ satisfying $j>k$ and $\{k,j\}\in E(G)$. Then, for any $1\leq i\leq s$, we have
$\mathrm{HS}_{i}(I^{s})=I^{s-i}\cdot\left(\frac{\alpha^{i}}{\prod_{k\in F}x_{\phi(k)}}\middle|\;F\subseteq[n-2]\text{\ and\ }|F|=i\right).$
@@
Let $e_{1},e_{2},\ldots,e_{n-2},e_{n-1}$ be all the edges of $G$. Since $\min(e_{1}),\ldots,\min(e_{n-1})$ are pairwise distinct, we may assume without loss of generality that $\min(e_{j})=j$ for every $j\in[n-1]$. It follows that $e_{j}=x_{j}x_{\phi(j)}$ for all $j\in[n-1]$. By [19, Lemmas 5.9-5.11], the edge ring $\mathbb{K}[x_{j}x_{\phi(j)}|1\leq j\leq n-1]$ has a trivial defining ideal. This implies that the quadratic monomials $e_{1},\ldots,e_{n-1}$ are algebraically independent over $\mathbb{K}$. Hence, every element of $\mathcal{G}(I(G)^{s})$ is uniquely written as
@@
where $\sum\limits_{j=1}^{n-1}s_{j}=s$ and each $s_{j}\geq 0$.
Take any $\beta=\frac{\alpha^{s-i}}{e_{1}^{a_{1}}\cdots e_{n-1}^{a_{n-1}}}\in\mathcal{G}(I^{s-i})$, where $e_{1}^{a_{1}}\cdots e_{n-1}^{a_{n-1}}\in\mathcal{G}(I(G)^{s-i})$, whence $\sum_{j=1}^{n-1}a_{j}=s-i$. For any subset $F\subseteq[n-2]$ with $|F|=i$, define
@@
where $b_{i}=1$ if $i\in F$, and $b_{i}=0$ otherwise. In particular, $b_{n-1}=0$. Note that $\gamma\in\mathcal{G}(I^{s})$, and $F\subseteq\operatorname{set}(\gamma)$ by the equation (3). This implies that ${\bf x}_{F}\gamma\in\mathrm{HS}_{i}(I^{s})$. Direct computation gives
$\beta\cdot\frac{\alpha^{i}}{\prod_{k\in F}x_{\phi(k)}}=\mathbf{x}_{F}\gamma\in\mathrm{HS}_{i}(I^{s}).$
Thus, the inclusion $\supseteq$ holds.
Conversely, take any $U\in\mathcal{G}(\mathrm{HS}_{i}(I^{s}))$. By definition of $\mathrm{HS}_{i}(I^{s})$, there exists a subset $F\subseteq[n]$ with $|F|=i$ and $V=\frac{\alpha^{s}}{e_{1}^{s_{1}}\cdots e_{n-1}^{s_{n-1}}}\in\mathcal{G}(I^{s})$ such that $U=\mathbf{x}_{F}V$ and $F\subseteq\mathrm{set}(V)$. By Equation (3), $F\subseteq[n-2]$ and $s_{k}>0$ for all $k\in F$. Define $c_{k}=s_{k}-1$ if $k\in F$ and $c_{k}=s_{k}$ otherwise; then $\sum_{k=1}^{n-1}c_{k}=s-i$, so $W=\frac{\alpha^{s-i}}{e_{1}^{c_{1}}\cdots e_{n-1}^{c_{n-1}}}\in\mathcal{G}(I^{s-i})$. Direct computation verifies
$W\cdot\frac{\alpha^{i}}{\prod_{k\in F}x_{\phi(k)}}=\mathbf{x}_{F}V=U,$
so $U\in I^{s-i}\cdot\left(\frac{\alpha^{i}}{\prod_{k\in F}x_{\phi(k)}}\middle|\;F\subseteq[n-2],|F|=i\right)$. Thus, the inclusion $\subseteq$ holds.
Combining both inclusions completes the proof.
@@
Let $G$ be a tree and let $I=I_{c}(G)$ be its complementary edge ideal. For $1\leq i\leq n-2$, the $i$-th homological shift algebra $\mathrm{HS}_{i}(\mathcal{R}(I))$ is generated exactly in degree $i$.
@@
By Theorem 5.2, we have $\mathrm{HS}_{i}(I^{s})=I^{s-i}\cdot\mathrm{HS}_{i}(I^{i})$ for all $s\geq i$. This inclusion implies $\mathrm{HS}_{i}(\mathcal{R}(I))$ is generated in degrees at most $i$. By Theorem 4.8, it follows that $\mathrm{HS}_{i}(I^{s})=0$ for all $s<i$ and no generators of $\mathrm{HS}_{i}(\mathcal{R}(I))$ can exist in degrees less than ii. Combining these two observations, this $\mathcal{R}(I)$-module is generated exactly in degree $i$.
The ideal $\left(\frac{\alpha^{i}}{\prod_{k\in F}x_{\phi(k)}}\middle|\;F\subseteq[n-2]\text{ and }|F|=i\right)$ is uniquely determined by $G$ and $i$ up to variable permutation, as it coincides with $\mathrm{HS}_{i}(I^{i})$. This implies the ideal is independent of the labeling of $G$, and we denote it by $J_{i}(G)$.
To study the properties of $J_{i}(G)$, we recall some foundational theory on polymatroidal ideals, following the notation from [6] and [21].
Given a monomial $u=x_{1}^{c_{1}}x_{2}^{c_{2}}\cdots x_{n}^{c_{n}}\in S$ where each $c_{i}\in\mathbb{Z}_{\geq 0}$, the $x_{i}$-degree of $u$, denoted $\deg_{x_{i}}(u)$, is defined as $\deg_{x_{i}}(u)=c_{i}$. An ideal $I\subset S$ generated by monomials of a single degree is called polymatroidal if it satisfies the exchange property, which states that for all $u,v\in\mathcal{G}(I)$ and every index $i$ with $\deg_{x_{i}}(u)>\deg_{x_{i}}(v)$, there exists an index $j$ such that $\deg_{x_{j}}(u)<\deg_{x_{j}}(v)$ and $x_{j}(u/x_{i})\in I$.
A stronger version, known as the strong exchange property, asserts that for all $u,v\in\mathcal{G}(I)$, every index $i$ with $\deg_{x_{i}}(u)>\deg_{x_{i}}(v)$, and all indices $j$ with $\deg_{x_{j}}(u)<\deg_{x_{j}}(v)$, the monomial $x_{j}(u/x_{i})$ lies in $I$. It is immediate that the strong exchange property implies the exchange property; consequently, every ideal satisfying the strong exchange property is polymatroidal.
A key class of polymatroidal ideals that satisfy the strong exchange property are ideals of Veronese type (also referred to as Veronese-type ideals): given a vector ${\bf a}=(a_{1},\ldots,a_{n})\in\mathbb{Z}_{\geq 0}^{n}$ and an integer $d$ with $|{\bf a}|=\sum_{i=1}^{n}a_{i}\geq d$, the ideal of Veronese type determined by $\bf a$} and $d$ is defined as
$I_{{\bf a},d}=\left({\bf x}^{b}\in S:|{\bf b}|=d,\,{\bf b}\leq{\bf a}\right),$
where ${\bf b}=(b_{1},\ldots,b_{n})\in\mathbb{Z}_{\geq 0}^{n}$, $|{\bf b}|=\sum_{i=1}^{n}b_{i}$ (the total degree of ${\bf x}^{b}$), and ${\bf b}\leq{\bf a}$ means $b_{i}\leq a_{i}$ for all $i$. Ideals of Veronese type have been studied in [10, 23, 22, 24]. By the work [22], it is known that ideals satisfying the strong exchange property are essentially ideals of Veronese type in the sense: For a monomial ideal $I\subset S$ generated in a single degree. Then $I$ satisfies the strong exchange property, if and only if, $I=(u)I_{{\bf a},d}$ for some monomial $u\in S$, and some ${\bf a}\in{\mathbb{Z}}_{\geq 0}^{n}$ and $d>0$.
@@
Let $G$ be a tree with $n$ vertices. Then $\mathrm{HS}_{i}(I_{c}(G)^{i})=J_{i}(G)$ is a polymatroidal ideal with the strong exchange property for each $i\in[n-2]$.
@@
Let $K_{i}(G)$ denote the monomial ideal of $S$ generated by all monomials of the form $\prod_{k\in F}x_{\phi(k)}$, where $F\subseteq[n-2]$ and $|F|=i$. For any $k\in[n-1]$, define
$b_{k}:=|\phi^{-1}(k)|=\left|\{j\in[n-2]:\phi(j)=k\}\right|.$
@@
$K_{i}(G)=\left(\prod\limits_{k=1}^{n-1}x_{k}^{a_{k}}\middle|\;\sum\limits_{k=1}^{n-1}a_{k}=i\text{\ and\ }0\leq a_{k}\leq b_{k}\ \text{for all}\ k\in[n-1]\right).$
Fix $i\in[n-2]$. For every $F\subseteq[n-2]$ such that $|F|=i$, set $a_{k}:=|F\cap\phi^{-1}(k)|$ for any $k\in[n-1]$. Then $0\leq a_{k}\leq b_{k}$ for all $k$, and
@@
Note that $\prod\limits_{k\in F}x_{\phi(k)}=\prod\limits_{k=1}^{n-1}x_{k}^{a_{k}}$, so the inclusion $\subseteq$ of the above claim holds.
Conversely, take any monomial $\prod\limits_{k=1}^{n-1}x_{k}^{a_{k}}$ such that $\sum\limits_{k=1}^{n-1}a_{k}=i$ and $0\leq a_{k}\leq b_{k}$ for all $k\in[n-1]$. Let $F=F_{1}\sqcup\cdots\sqcup F_{n-1}$, where $F_{k}$ is a subset of $a_{k}$ elements in $\phi^{-1}(k)$, and by convention, $F_{k}=\emptyset$ if $a_{k}=0$. Then $\prod_{k\in F}x_{\phi(k)}=\prod\limits_{k=1}^{n-1}x_{k}^{a_{k}}$, so the inclusion $\supseteq$ holds. This proves the claim. Therefore, $K_{i}(G)$ coincides with the Veronese-type ideal $I_{\mathbf{b},i}$, where ${\bf b}=(b_{1},\ldots,b_{n-1}).$
By putting $b_{n}=0$, we conclude that:
$J_{i}(G)=\left(\prod\limits_{k=1}^{n}x_{k}^{i-a_{k}}\middle|\;\sum\limits_{k=1}^{n-1}a_{k}=i\text{\ and\ }0\leq a_{k}\leq b_{k}\ \text{for all}\ k\in[n]\right)$
$=\left(\prod\limits_{k=1}^{n}x_{k}^{c_{k}}\middle|\;\sum\limits_{k=1}^{n-1}c_{k}=(n-2)i\text{\ and\ }i-b_{k}\leq c_{k}\leq i\ \text{for all}\ k\in[n]\right)$
$=\left(\prod\limits_{k=1}^{n}x_{k}^{i-b_{k}}\right)\left(\prod\limits_{k=1}^{n-1}x_{k}^{c_{k}-i+b_{k}}\middle|\;\sum\limits_{k=1}^{n-1}c_{k}=(n-2)i\text{\ and\ }i-b_{k}\leq c_{k}\leq i\ \text{for all}\ k\in[n]\right)\$
$=\left(\prod\limits_{k=1}^{n}x_{k}^{i-b_{k}}\right)\left(\prod\limits_{k=1}^{n-1}x_{k}^{a_{k}}\middle|\;\sum\limits_{k=1}^{n-1}a_{k}=n-2-i\text{\ and\ }0\leq a_{k}\leq b_{k}\ \text{for all}\ k\in[n]\right)$
@@
Here, the second to last equality holds because of $\sum_{k=1}^{n-1}b_{k}=n-2$. In the last equality, we use the convention that $K_{0}(G)=S$. Therefore, $J_{i}(G)$ is a polymatroidal ideal with the strong exchange property.
Let $I$ be a polymatroidal ideal. A cage for $I$ is an integral vector $\mathbf{c}\in\mathbb{N}^{n}$ such that $\mathbf{u}(i)\leq\mathbf{c}(i)$ for all $\mathbf{u}$ with $\mathbf{x}^{\mathbf{u}}\in\mathcal{G}(I)$ and all $i\in[n]$. Given a cage $\mathbf{c}$ of $I$, the ideal generated by $\mathbf{x}^{\mathbf{c}-\mathbf{u}}$ for all $\mathbf{u}\in\mathcal{G}(I)$ is itself a polymatroidal ideal; we refer to this object as the dual polymatroidal ideal of $I$ with respect to $\mathbf{c}$. For more details, see [6, 10]. By virtue of this definition, $J_{i}(G)$ is precisely the dual polymatroidal ideal of $K_{i}(G)$ with respect to the $n$-tuple cage $\mathbf{c}=(i,\dots,i)$.
In Proposition 5.4, we show that the $i$-th homological shift ideal $\mathrm{HS}_{i}(I_{c}(G)^{i})$ (denoted $J_{i}(G)$) is a Veronese-type ideal after dividing by a suitable monomial. In fact, every Veronese-type ideal arises in this manner.
Recall that a caterpillar tree is a tree $T$ containing a path $P$ such that every vertex of $T$ is either on $P$ or adjacent to a vertex of $P$. Caterpillar trees were first investigated by Harary and Schwenk [18]. A leaf edge is an edge connecting a leaf vertex (a vertex of degree 1) to another vertex in the graph.
@@
Given a vector $\mathbf{b}=(b_{1},\ldots,b_{n})\in\mathbb{Z}_{>0}^{n}$ and a positive integer $d\leq|\mathbf{b}|$, where $|\mathbf{b}|=\sum_{j=1}^{n}b_{j}$. Set $\sigma_{k}=\sum_{j=1}^{k}b_{j}$ for all $k\in[n]$, and let $i=|\mathbf{b}|-d$. Then there exists a caterpillar tree $T$ with $|\mathbf{b}|+2$ vertices such that
$\mathrm{HS}_{i}\bigl(I_{c}(T)^{i}\bigr)=uI,$
where $u$ is a monomial, and $I\subseteq\mathbb{K}[x_{j}\mid 1\leq j\leq|\mathbf{b}|+2]$ is a Veronese-type ideal generated by all monomials of the form $x_{\sigma_{1}+1}^{a_{1}}x_{\sigma_{2}+1}^{a_{2}}\cdots x_{\sigma_{n}+1}^{a_{n}}$ with $0\leq a_{j}\leq b_{j}$ for all $j\in[n]$ and $\sum_{t=1}^{n}a_{t}=d$.
We remark that the ideal $I$ is essentially identical to the ideal $I_{\mathbf{b},d}\subset S$, despite their residing in the distinct ambient rings.